# Efficient Swept Volume-Based Trajectory Generation for Arbitrary-Shaped Ground Robot Navigation

### v1

@@ VII-A1 Baselines
We evaluate the proposed method against two state-of-the-art any-shape robot motion planners: SVSDF [6] and RC-ESDF [4]. Both baselines deploy A* as their front-end global path search but differ fundamentally in collision-free orientation storing. SVSDF extends A* to explore robot kernel position and orientation during node expansions, storing the full SE(2) state at each node. In contrast, RC-ESDF adopts traditional A* without considering orientation in this stage, parameterizing trajectories as uniform B-splines. Neither baseline inherently supports topological diversity paths. For fairness, we unify all methods using MINCO [17] for trajectory representation and generate reference topological paths via Zhou’s method [15]. RC-ESDF is guided by $\mathbb{R}^{2}$ paths with collision avoidance. At the same time, SVSDF initializes SE(2) optimization by discretizing topological paths and solving for feasible orientations at each point.
@@ VII-B Real-world Tests
We validate our method in a real-world indoor office environment using a robot modified with a goods-shelf structure to emulate non-convex geometry and customer drink delivery capability. The platform employs a Livox Mid360 LiDAR (with built-in IMU) and an Intel NUC onboard computer (i7-1360P CPU). High-accuracy and robust localization and mapping modules (e.g., [18, 19, 20]) are necessary for the robot to perform precise obstacle avoidance. We deploy FAST-LIO2 [20] in this work, enabling aggressive maneuvers in narrow spaces. Experimental results (Table I) demonstrate the method’s efficiency.

### v2

@@ VII-A1 Baselines
We evaluate the proposed method against two state-of-the-art any-shape robot motion planners: SVSDF [SVSDF] and RC-ESDF [geng]. Both baselines deploy A* as their front-end global path search but differ fundamentally in collision-free orientation storing. SVSDF extends A* to explore robot kernel position and orientation during node expansions, storing the full SE(2) state at each node. In contrast, RC-ESDF adopts traditional A* without considering orientation in this stage, parameterizing trajectories as uniform B-splines. Neither baseline inherently supports topological diversity paths. For fairness, we unify all methods using MINCO [MINCO] for trajectory representation and generate reference topological paths via Zhou’s method [topology]. RC-ESDF is guided by $\mathbb{R}^{2}$ paths with collision avoidance. At the same time, SVSDF initializes SE(2) optimization by discretizing topological paths and solving for feasible orientations at each point.
@@ VII-B Real-world Tests
We validate our method in a real-world indoor office environment using a robot modified with a goods-shelf structure to emulate non-convex geometry and customer drink delivery capability. The platform employs a Livox Mid360 LiDAR (with built-in IMU) and an Intel NUC onboard computer (i7-1360P CPU). High-accuracy and robust localization and mapping modules (e.g., [fastlivo1, fastlivo2, FASTLIO2]) are necessary for the robot to perform precise obstacle avoidance. We deploy FAST-LIO2 [FASTLIO2] in this work, enabling aggressive maneuvers in narrow spaces. Experimental results (Table I) demonstrate the method’s efficiency.